# A Lightweight Method for Defending Against UAF Vulnerabilities

### v1

@@ 6. Limitations and Future Work
We have demonstrated through experiments that LightDE is very effective at eliminating dangling pointers. However, LightDE still faces some limitations. The first limitation of LightDE comes from the imprecision of static object abstraction. As analyzed in our experiments, the reason of the high memory overhead in projects 471.omnetpp and 483.xalancbmk is the inaccuracy of the pointer analysis results. This inaccuracy is due to the fact that the object abstraction method we use in the pointer analysis is too coarse. In our pointer analysis, we abstract the memory allocation point as a single object. In a static case, when a pointer at offset $f$ of a static object $o1$ points to another static object $o2$, the corresponding runtime case is that all runtime objects created at the allocation point of $o1$ are considered to have pointers at offset $f$ that point to all runtime objects created at the allocation point of $o2$. This leads to checking too many irrelevant pointers when eliminating heap dangling pointers. This excessive checking of unrelated pointers consumes a lot of time during the elimination process, causing the LightDE thread to be unable to promptly release the objects passed by the original program thread through the lock-free queue, resulting in very high memory overhead. In our future work, we will investigate more accurate object abstraction methods to avoid excessive checks of irrelevant heap pointers during the elimination of dangling pointers.
LightDE has the issue of false positives. In some cases, programmers deliberately use dangling pointers but not for memory access. For example, some programs may use dangling pointers to calculate the address of another memory location, or convert dangling pointers to integers to compare numerical values with other integers. For instance, suppose we have two dangling pointers, $p$ and $q$, which are converted to integers and then compared using $p>q$. Since LightDE sets dangling pointers to null pointers, the comparison results change, leading to false positives. In this case, LightDE incorrectly sets the dangling pointer to null. This issue is a common problem faced by all methods that defend against UAF vulnerabilities by eliminating dangling pointers.
LightDE also has the issue of false negatives. Our method only eliminates dangling pointers in memory and cannot eliminate those in registers. When compiling a program, the compiler promotes local pointers, such as stack pointers and function parameters, to registers to improve the program’s execution speed. During program execution, register pointers reside in registers while functions are active, and when functions are inactive, register pointers are temporarily stored on the stack. Therefore, when memory is released, we cannot determine the memory space of the register pointer. Therefore, register pointers cannot be eliminated by LightDE, potentially leading to false negatives. To our knowledge, this is a common limitation of all methods that defend against UAF vulnerabilities by eliminating dangling pointers, not just ours.
7. RELATED WORK
There are two main methods for eliminating dangling pointers: The first method uses taint tracking technology to track pointers and eliminate them after the heap memory is freed. The second method involves storing pointer addresses in a specific global structure and then eliminating these pointers after the heap memory is freed. Undangle (Caballero et al., 2012) allocates a taint identifier for pointers during memory allocation and dynamically tracks the propagation of taints. When memory is freed, Undangle identifies dangling pointers through the taint identifier. This method can eliminate all dangling pointers, whether they are in memory, in registers, or have been converted to integers for propagation. However, this method incurs extremely high runtime and memory overhead, making it impractical for real-world applications and suitable only for vulnerability detection.
@@ 7. RELATED WORK
Another method is to store the pointer addresses in a carefully designed global structure and eliminate the corresponding dangling pointers when heap memory is freed. This method is adopted by DangNULL (Lee et al., 2015), FreeSentry (Younan, 2015), DangSan (Van Der Kouwe et al., 2017), HeapExpo (Shen and Dolan-Gavitt, 2020), and pSweeper (Liu et al., 2018). DangNULL and FreeSentry insert a specific function after pointer assignment instructions during compilation. This function is used to look up the objects pointed to by the pointers and add the pointers to the object’s pointer set. DangNULL and FreeSentry maintain the object pointed to by the pointer in real-time. When the object pointed to by a pointer changes, DangNULL and FreeSentry remove the pointer from the pointer set of the previously pointed-to object and add it to the pointer set of the newly pointed-to object. DangNULL uses a tree structure to store the object’s pointer set, while FreeSentry uses a two-level mapping table. DangSan uses a system log-like method to record the objects pointed to by pointers. DangSan inserts a function after pointer assignment instructions to add the pointer to the pointer set of the object it points to. If the target of a pointer changes, it does not remove the pointer from the set of the previously pointed-to object. This approach avoids the real-time maintenance of object pointer sets, resulting in lower performance overhead. When an object is freed, DangSan checks whether the pointers in the object’s pointer set still point to the freed object. HeapExpo uses the same method as DangSan but prohibits the compiler from optimizing local pointers into registers, resulting in very high performance overhead. To further improve performance, pSweeper adopts an asynchronous method to eliminate dangling pointers. During program execution, pSweeper only stores pointers in its carefully designed global data structure, avoiding the need to look up the objects pointed to by the pointers. pSweeper runs a separate thread on an idle CPU core. This thread continuously scans the pointers recorded in the global data structure to check if they point to freed memory and eliminates dangling pointers. However, all these methods require interrupting the program’s execution flow to store the memory addresses of pointers in their global data structures and to look up the objects pointed to by the pointers.
8. Conclusion
@@
This paper proposes LightDE, a lightweight method for eliminating dangling pointers to defend against UAF vulnerabilities. LightDE stores pointer points-to information in the program’s data segment at compile time based on our pointer analysis results, avoiding the need to store pointer memory addresses and look up pointer targets during program execution. To achieve this, we present a structure-sensitive pointer analysis method for analyzing the type information of objects. Additionally, we introduce a method for mapping static pointer analysis results to runtime information to accurately eliminate dangling pointers. We implemented a prototype system of LightDE and validated its effectiveness and efficiency.

### v4

@@ VI Limitations and Future Work
We have demonstrated through experiments that LightDE is very effective at eliminating dangling pointers. However, LightDE still has a major limitation. The limitation of LightDE arises from the imprecision of static object abstraction. This inaccuracy is due to the fact that the object abstraction method we use in the pointer analysis is too coarse. In our pointer analysis, we abstract the memory allocation point as a single object. In a static case, when a pointer at offset $f$ of static object $o1$ points to another static object $o2$, the corresponding runtime case is that all runtime objects created at the allocation point of $o1$ are considered to have pointers at offset $f$ that point to all runtime objects created at the allocation point of $o2$. This leads to checking too many irrelevant pointers when eliminating heap dangling pointers. This excessive checking of irrelevant pointers consumes a lot of time during the elimination process, causing the LightDE thread to be unable to timely release objects that are pending release, which are passed through the lock-free queue by the original program’s thread, resulting in very high memory overhead. In our future work, we will investigate more accurate object abstraction methods.
VII RELATED WORK
@@
Existing methods to defend against UAF vulnerabilities by eliminating dangling pointers include DangNULL [22], FreeSentry [7], DangSan [23], HeapExpo [25], pSweeper [24], and ISDE [9]. DangNULL and FreeSentry insert a specific function after pointer assignment instructions during compilation. This function is used to look up the objects pointed to by the pointers and add the pointers to the object’s pointer set. DangNULL and FreeSentry maintain the objects pointed to by the pointers in real time. When the object pointed to by a pointer changes, DangNULL and FreeSentry remove the pointer from the pointer set of the previously pointed-to object and add it to the pointer set of the newly pointed-to object. DangNULL uses a tree structure to store the object’s pointer set, while FreeSentry uses a two-level mapping table. DangSan uses a method similar to a system log to record pointers. DangSan inserts a function after pointer assignment instructions to add the pointer to the pointer set of the object it points to. If the target of a pointer changes, it does not remove the pointer from the set of the previously pointed-to object. This approach avoids the real-time maintenance of object pointer sets, resulting in lower performance overhead. When an object is freed, DangSan checks whether the pointers in the object’s pointer set still point to the object. HeapExpo uses the same method as DangSan but prohibits the compiler from optimizing local pointers into registers, resulting in very high performance overhead. To further improve performance, pSweeper adopts an asynchronous method to eliminate dangling pointers. During program execution, pSweeper only stores pointers in its carefully designed global data structure, avoiding the need to look up the objects pointed to by the pointers. pSweeper runs a separate thread on an idle CPU core. This thread continuously scans the pointers recorded in the global data structure to check if they point to freed memory and eliminates dangling pointers. However, all of these methods require recording the memory addresses of pointers at runtime, and DangNULL, FreeSentry, DangSan, and HeapExpo also need to locate the objects pointed to by the pointers during program execution. Finally, it is worth noting that the previous method, ISDE, also attempted to eliminate dangling pointers by combining pointer analysis techniques [9]. However, ISDE mainly focused on eliminating dangling pointers in registers. ISDE uses a flow-sensitive pointer analysis method to determine the targets of pointers. However, the flow-sensitive pointer analysis method can only analyze the objects pointed to by pointers in single-threaded applications. Therefore, ISDE can only eliminate dangling pointers in single-threaded programs.
VIII Conclusion
@@
This paper proposes LightDE, a lightweight method for eliminating dangling pointers to defend against UAF vulnerabilities. LightDE stores pointer points-to information in the program’s data segment at compile time based on our pointer analysis results, avoiding the need to store pointer memory addresses and look up pointer targets during program execution. To achieve this, we present a structure-sensitive pointer analysis method for analyzing the points-to information and the type information of objects. We implemented a prototype system of LightDE and validated its effectiveness and efficiency.